# Probing $p$-wave effects in spin-density separation of Bose mixtures by dynamic structure factor

### v1

@@ I Introduction
At present, there is considerable and sustained interest in delving into the realm of $p$-wave pairing-dominated physics within the framework of quantum gas mixtures, distinguished by an orbital angular momentum of $l=1$ [1]. The motivation behind these interests is twofold. First, compared to $s$-wave pairing-dominated physics, pairing with finite angular momentum hosts richer structural complexity and plays a critical role in novel superconductors and superfluids [2]. This is exemplified by two landmark systems: (i) superfluid 3He where the $p$-wave pairing governs the symmetry-dependent emergence of gapless (A-phase) and gapped (B-phase) superfluidity [3], and (ii) Sr2RuO4, a proposed two-dimensional chiral $p_{x}+ip_{y}$ superconductor harboring Majorana edge modes with non-Abelian statistics—a cornerstone for topological quantum computing [4, 5]. Second, advances in experimental techniques—notably Feshbach resonance [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]—enable precise tuning of $p$-wave interactions in ultracold gases. This unprecedented control over inter-intraspecies interactions revitalizes efforts to realize $p$-wave superfluids while bridging insights to condensed-matter systems, such as unconventional superconductors [5, 16].
Along this research line of pursuing novel quantum phases in quantum gas mixtures, a paramount prerequisite is the precise control of miscibility, governed by the interplay between atomic interactions and experimental geometry [17, 18]. For $s$-wave systems, miscibility in binary Bose-Einstein condensates (BECs) is well-characterized across static [19, 20, 21] and dynamical regimes [22, 23]. In contrast, $p$-wave interactions introduce unique paradigms: mean-field studies predict their ability to regulate phase separation [24], while quantum corrections stabilize exotic states (e.g., polar droplets) and extend thermodynamics to higher partial waves [25, 26]. However, critical gaps persist in understanding the interplay between tunable $p$-wave and $s$-wave interactions—particularly their combined effects on non-equilibrium dynamics and experimental observables—posing both a theoretical and technical frontier.
@@ II.1 Partition function of the system
$\mathcal{L}\left[\bar{\Psi},\Psi\right]$ $=$ $\bar{\Psi}\left({\bf r},\tau\right)\left[\hbar\partial_{\tau}-\frac{\hbar^{2}}{2m}\nabla^{2}-\mu\right]\Psi\left({\bf r},\tau\right)-g_{p}\big{|}\Psi_{1}\left({\bf r},\tau\right)\nabla\Psi_{2}\left({\bf r},\tau\right)-\Psi_{2}\left({\bf r},\tau\right)\nabla\Psi_{1}\left({\bf r},\tau\right)\big{|}^{2}$ (2)
$+$ $\frac{g+g_{12}}{4}\left(\left|\Psi_{1}\left({\bf r},\tau\right)\right|^{2}+\left|\Psi_{2}\left({\bf r},\tau\right)\right|^{2}\right)+\frac{g-g_{12}}{4}\left(\left|\Psi_{1}\left({\bf r},\tau\right)\right|^{2}-\left|\Psi_{2}\left({\bf r},\tau\right)\right|^{2}\right),$
@@
Table 1: Ground state phase diagram of the two-component Bose mixture under variation of the coefficient $y$ and $z$.
@@
$z>0$ Quantum Droplet under Ref. [26]
@@
$10<y<$ 1, $z=0$ Spin-density seperation Ref. [31]
$00<y<1,z>$ 0 Spin-Density Separation This work
@@
$y>1,z>0$ Phase Seperation under Ref. [24]
@@
In the case where $z=0$, corresponding to the absence of $p$-wave interaction, (i) a quantum droplet phase exists for $y<-1$, as first studied in Ref. [36]; (ii) for $-1<y<1$, the system remains miscible, and within the range $10<y<$ 1, spin-density separation can occur, enabling further experimental exploration of interaction effects on the system [24, 31]; (iii) when $y>1$, the system exhibits phase separation [44].
Upon introducing $p$-wave interactions, tuning $z$ gives rise to distinct physical phenomena in the two-component system, such as $p$-wave-affected quantum droplets [26], $p$-wave-affected phase separation [24], and $p$-wave-affected spin-density separation, which is the focus of this work. These effects further enrich the phase behavior of the system and open up new avenues for theoretical and experimental investigation. In Sec. IV, we provide a comprehensive discussion of $p$-wave affected spin-density separation and present the corresponding theoretical framework for probing this phenomenon.
II.2 Bogliubov excitations
@@
$S_{2}\left[\eta_{1}^{*},\eta_{1},\eta_{2}^{*},\eta_{2}\right]$ $=$ $-\frac{\hbar}{2}\sum_{\bm{k}\neq 0,n}{\bf\Phi}_{\bm{k}n}\cdot\bm{G}_{\bm{k}n}^{-1}\cdot{\bf\Phi}_{\bm{k}n}^{\dagger},$ (4)
@@
Figure 1: (a) Dimensionless excitation spectrum $f_{+}(k^{\prime})$ in Eq. (19) for different values of $z=g_{3}n_{0}$ from $0-2$. (b) Dimensionless excitation spectrum $f_{-}(k^{\prime})$ in Eq. (19) for different values of $z=g_{3}n_{0}$ from $0-2$. Here the dimensionless interation parameters $y=0.2$.
@@ III Equation of state
In Eq. (21), the first and second terms on the right-hand side represent the mean-field contribution, whereas all subsequent terms account for the beyond-mean-field corrections stemming from quantum fluctuations. It is noteworthy that the last three terms in Eq. (21) are incorporated to mitigate ultraviolet divergence through an appropriate renormalization of the coupling constant [53]. In the continuum limit, the summation in Eq. (21) can be systematically replaced by an integral, yielding the analytical expression for the ground-state energy of the model system as follows:
@@
IV PROBING SPIN-DENSITY SEPERATION USING BRAGG SPECTROSCOPY
@@
Figure 3: Dimensionless dynamic structure factor function $J(\omega^{\prime})$ in the hydrodynamic regime and at zero temperature for $y=0.5$ and $k=1.25$ shows diffrent distinct peaks from $z=0-1$ corresponding to the density and the spin waves, centered at $1$ and $\beta$.
@@
Here, the ground-state compressibility $\chi_{\rho(\sigma)}$ is related to the ground-state energy $E$ through the expression $\chi^{-1}=\frac{1}{V}\frac{\partial^{2}E}{\partial n^{2}}$, where $V$ denotes the constant system volume and $n=N/V$ represents the density. The compressibility for the density degree of freedom is given by $\chi_{\rho}^{-1}=g_{\rho}+2g_{\rho}^{5/2}n_{0}^{1/2}m_{\rho}^{3/2}/(\pi^{2}\hbar^{3})$, and the compressibility for the spin-density degree of freedom is given by $\chi^{-1}_{\sigma}=g_{\sigma}+\frac{2g_{\sigma}^{5/2}m^{3/2}n_{0}^{1/2}z^{2}}{\pi^{2}\hbar^{3}(z+1)^{7/2}}-\frac{4g_{\sigma}^{5/2}m^{3/2}n_{0}^{1/2}z}{\pi^{2}\hbar^{3}(z+1)^{5/2}}+\frac{2g_{\sigma}^{5/2}m^{3/2}n_{0}^{1/2}}{\pi^{2}\hbar^{3}(z+1)^{3/2}}$. The sound velocity is similarly determined as $v_{\rho\left(\sigma\right)}=\left(\frac{V}{mn}\frac{\partial^{2}E}{\partial V^{2}}\right)^{1/2}$ with a constant particle number $N$, expressed for the density degree of freedom as $\nu_{\rho}=\left(g_{\rho}n_{0}/m_{\rho}+2g_{\rho}^{5/2}m_{\rho}^{1/2}n_{0}^{3/2}/\left(\pi^{2}\hbar^{3}\right)\right)^{1/2}$, and for the spin-density degree of freedom as $\nu_{\sigma}=\left(\frac{2g_{\sigma}^{5/2}m^{1/2}n_{0}^{3/2}z^{2}}{\pi^{2}\hbar^{3}\left(z+1\right)^{5/2}}+\frac{2g_{\sigma}^{5/2}m^{1/2}n_{0}^{3/2}}{\pi^{2}\hbar^{3}\left(z+1\right)^{1/2}}-\frac{4g_{\sigma}^{5/2}m^{1/2}n_{0}^{3/2}z}{\pi^{2}\hbar^{3}\left(z+1\right)^{3/2}}+\frac{g_{\sigma}n_{0}\left(1+z\right)}{m}\right)^{1/2}$. Additionally, in calculating the damping rates, we approximate that the density fluctuations in the $p$-wave interaction vertex term predominantly couple to the spin-wave degrees of freedom, consistent with our prior analysis. This enables us to apply $m_{\rho}=m$ and $m_{\sigma}=m/\left(1+z\right)$ to the damping rates of the density and spin modes, respectively. Following Ref. [57], the decay rates for the density and spin-density modes are given by $\Gamma_{\rho\left(\sigma\right)}=3\hbar k^{5}/\left(640\pi m_{\rho\left(\sigma\right)}n_{0}\right)$. Using these decay rates, we can reexpress the DSFs as follows:
$S_{\rho\left(\sigma\right)}\left(\omega\right)$ $=$ $\chi_{\rho}J_{\rho\left(\sigma\right)}\left(\omega^{\prime}\right)$ (42)
@@
$J_{\rho}\left(\omega^{\prime}\right)$ $=$ $\frac{\Gamma_{\rho}/v_{\rho}k}{\left(\omega^{\prime}-1\right)^{2}+\left(\Gamma_{\rho}/v_{\rho}k\right)^{2}},$ (43)
@@
$J_{\sigma}\left(\omega^{\prime}\right)$ $=$ $\frac{\alpha\Gamma_{\sigma}/v_{\rho}k}{\left(\omega^{\prime}-\beta\right)^{2}+\left(\Gamma_{\sigma}/v_{\rho}k\right)^{2}}.$ (44)
Here, $\omega^{\prime}=\omega/v_{\rho}k$ represents the dimensionless frequency, with additional dimensionless coefficients defined as $\alpha=\chi_{\sigma}v_{\sigma}/\chi_{\rho}v_{\rho}$ and $\beta=v_{\sigma}/v_{\rho}$. Setting $k=1.25$ and $y=0.5$, the density compressibility $\chi_{\rho}$ can be approximately regarded as constant. The dimensionless DSF functions $J_{\rho\left(\sigma\right)}$ reflect the variation with frequency $\omega^{\prime}$ under different values of $z$, as depicted in Fig. 3. From Fig. 3, we observe that the density wave typically exhibits a higher sound velocity and a broader peak due to Beliaev damping, whereas the spin wave tends to display a narrower and more distinct response. Notably, the peak of the spin-density wave progressively shifts toward higher frequencies as the strength of the $p$-wave interaction increases. This offers a clear visualization of the influence of $p$-wave interactions on the system, providing a direct experimental signature detectable through Bragg scattering experiments.
@@ V CONCLUSION
The primary contribution of this article lies in deriving the EOS under $p$-wave interactions and elucidating their influence on a two-component Bose-Bose mixture through spin-density separation. Beyond providing explicit expressions for the LHY-order corrections to the ground-state energy and quantum depletion, as detailed in Sec. III, this work offers a visualization scheme via the DSFs and a comprehensive analysis of the theoretical results, as presented in Sec. IV. A key finding is that $p$-wave interactions, parameterized by $z$, induce a tunable splitting in the DSFs, reflecting affected particle-number fluctuations in the spin degrees of freedom, which distinguishes this system from purely $s$-wave-dominated mixtures. Furthermore, the validity of the Bogoliubov approximation employed in our calculations is retrospectively confirmed by estimating the quantum depletion $\left(N-N_{0}\right)/N$, utilizing Fig. 2 (b). In typical Bose-Bose mixture experiments, with parameters $n\approx 3\times 10^{14}\leavevmode\nobreak\ \text{cm}^{-3}$, $a\approx 2.7517\leavevmode\nobreak\ \text{nm}$ [58], the quantum depletion in Eq. (26) is evaluated as $\left(N-N_{0}\right)/N\approx 0.0028\times I\left(z\right)$ shown in Fig. 2 (b), affirming the applicability of the Bogoliubov approximation [59]. The experimental realization of our scenario relies on precise control of three key parameters: the $s$-wave inter- and intra-species scattering lengths ($a_{12}$ and $a$) and the $p$-wave scattering volume ($v_{p}$), all of which are highly tunable via Feshbach resonance [54, 15]. We anticipate that these predicted features, particularly the $p$-wave-induced spin-density separation observable through Bragg spectroscopy, will be validated in future experiments.
@@
In summary, this paper examines the effects of $p$-wave interactions on the EOS and spin-density separation within a Bose-Bose mixture. At the Gaussian order, we demonstrate that $p$-wave interactions fully decouple the spin-density degree of freedom, resulting in a compression effect confined to its effective mass, as evidenced by the renormalization captured in Sec. IV. This decoupling, driven by the momentum-dependent nature of $p$-wave interactions, not only manifests as distinct DSF peaks but also enriches the system’s phase behavior beyond traditional $s$-wave frameworks. By bridging theoretical predictions with experimental feasibility, this study enhances our understanding of interaction-driven phenomena in multi-component Bose systems and provides robust theoretical guidance for future explorations in higher-dimensional quantum gases, potentially advancing the study of exotic quantum states influenced by $p$-wave interactions.
ACKNOWLEDGMENTS

### v3

# Probing $p$-wave effects in spin-density separation of Bose mixtures with the dynamic structure factor

@@ I Introduction
At present, there is considerable and sustained interest in delving into the realm of $p$-wave pairing-dominated physics within the framework of quantum gas mixtures, distinguished by an orbital angular momentum of $l=1$ [1]. The motivation behind these interests is twofold. First, compared to $s$-wave pairing-dominated physics, pairing with finite angular momentum hosts richer structural complexity and plays a critical role in novel superconductors and superfluids [2]. This is exemplified by two landmark systems: (i) superfluid 3He where the $p$-wave pairing governs the symmetry-dependent emergence of gapless (A-phase) and gapped (B-phase) superfluidity [3], and (ii) Sr2RuO4, a proposed two-dimensional chiral $p_{x}+ip_{y}$ superconductor harboring Majorana edge modes with non-Abelian statistics—a cornerstone for topological quantum computing [4, 5]. Second, advances in experimental techniques, notably Feshbach resonance [6, 7, 8, 9, 10, 11, 12, 13, 14, 15], enable precise tuning of $p$-wave interactions in ultracold gases. This unprecedented control over inter-intraspecies interactions revitalizes efforts to realize $p$-wave superfluids while bridging insights to condensed-matter systems, such as unconventional superconductors [5, 16].
Along this research line of pursuing novel quantum phases in quantum gas mixtures, a paramount prerequisite is the precise control of miscibility, governed by the interplay between atomic interactions and experimental geometry [17, 18]. For $s$-wave systems, miscibility in binary Bose-Einstein condensates (BECs) is well-characterized across static [19, 20, 21] and dynamical regimes [22, 23]. In contrast, $p$-wave interactions introduce unique paradigms: mean-field studies predict their ability to regulate phase separation [24], and notably, they can dramatically increase the effective mass of polaronic neutrons, causing it to diverge as the alpha condensate density approaches a critical threshold [25]. Meanwhile, quantum corrections stabilize exotic states (e.g., polar droplets) and extend thermodynamics to higher partial waves [26, 27]. However, critical gaps persist in understanding the interplay between tunable $p$-wave and $s$-wave interactions—particularly their combined effects on non-equilibrium dynamics and experimental observables—posing both a theoretical and technical frontier.
@@ II.1 Partition function of the system
$\mathcal{L}\left[\bar{\Psi},\Psi\right]$ $=\bar{\Psi}\left({\bf r},\tau\right)\left[\hbar\partial_{\tau}-\frac{\hbar^{2}}{2m}\nabla^{2}-\mu\right]\Psi\left({\bf r},\tau\right)-g_{p}\left|\Psi_{1}\left({\bf r},\tau\right)\nabla\Psi_{2}\left({\bf r},\tau\right)-\Psi_{2}\left({\bf r},\tau\right)\nabla\Psi_{1}\left({\bf r},\tau\right)\right|^{2}$
$\quad+\frac{g+g_{12}}{4}\left(\left|\Psi_{1}\left({\bf r},\tau\right)\right|^{2}+\left|\Psi_{2}\left({\bf r},\tau\right)\right|^{2}\right)^{2}+\frac{g-g_{12}}{4}\left(\left|\Psi_{1}\left({\bf r},\tau\right)\right|^{2}-\left|\Psi_{2}\left({\bf r},\tau\right)\right|^{2}\right)^{2}.$ (2)
@@
Table 1: Ground state phase diagram of the two-component Bose mixture under variation of the coefficient $y$ and $z$,where negative parameters represent attractive interactions, while positive parameters indicate repulsive interactions.
@@
y=0, $z<0$ and $z\geq 0$ Quantum droplet under Ref. [27]
@@
$10<y<$ 1, $z=0$ Spin-density separation Ref. [32]
$0<y<1,z\geq 0$ Spin-density separation This work
@@
$y>1,z>0$ Phase separation under Ref. [24]
@@
In the case where $z=0$, corresponding to the absence of $p$-wave interaction, (i) a quantum droplet phase exists for $y<-1$, as first studied in Ref. [37]; (ii) for $-1<y<1$, the system remains miscible, and within the range $10<y<$ 1, spin-density separation can occur, enabling further experimental exploration of interaction effects on the system [24, 32]; (iii) when $y>0$, a mixed bubble phase emerges as discussed in Ref. [45]; (iv) when $y>1$, the system exhibits phase separation [48].
Upon introducing $p$-wave interactions, tuning $z$ gives rise to distinct physical phenomena in the two-component system, such as $p$-wave-affected quantum droplets [27], $p$-wave-affected phase separation [24], $p$-wave-affected magnetoroton [46, 47], and $p$-wave-affected spin-density separation, which is the focus of this work. These effects further enrich the phase behavior of the system and open up new avenues for theoretical and experimental investigation. In Sec. IV, we provide a comprehensive discussion of $p$-wave affected spin-density separation and present the corresponding theoretical framework for probing this phenomenon.
II.2 Bogoliubov excitations
@@
$S_{2}\left[\eta_{1}^{*},\eta_{1},\eta_{2}^{*},\eta_{2}\right]$ $=$ $-\frac{\hbar}{2}\sum_{\bm{k}\neq 0,n}{\bf\Phi}_{\bm{k}n}\bm{G}_{\bm{k}n}^{-1}{\bf\Phi}_{\bm{k}n}^{\dagger},$ (4)
@@
Figure 1: (a) Dimensionless excitation spectrum $f_{+}(k^{\prime})$ in Eq. (19) for different values of $z=g_{3}n_{0}$ from $0-2$. (b) Dimensionless excitation spectrum $f_{-}(k^{\prime})$ in Eq. (19) for different values of $z=g_{3}n_{0}$ from $0-2$. Here the dimensionless interaction parameter $y=0.2$.
@@ III Equation of state
In Eq. (21), the first and second terms on the right-hand side represent the mean-field contribution, whereas all subsequent terms account for the beyond-mean-field corrections stemming from quantum fluctuations. It is noteworthy that the last three terms in Eq. (21) are introduced to remove power-law ultraviolet divergences in the momentum summation, a process that can be absorbed into an appropriate renormalization of the coupling constant $g_{p}$ [57]. While this approach aligns with established techniques for handling divergences, as exemplified by $s$-wave renormalization in Ref. [58], we refrain from an explicit derivation of the renormalized $g_{p}$ here. Instead, this subtraction ensures a finite LHY correction, consistent with the physical objectives of our study; the detailed derivation of these three terms is provided in Appendix B. When we take the limit $z=0$ or further set $y=0$, Eq. (21) reduces to the form presented in Refs. [38, 55], double-checking the reasonableness and correctness of our approach. In the continuum limit, the summation in Eq. (21) can be systematically replaced by an integral, yielding the analytical expression for the ground-state energy of the model system as follows:
@@
IV PROBING SPIN-DENSITY SEPARATION USING BRAGG SPECTROSCOPY
@@
Figure 3: Dimensionless dynamic structure factor function $J(\omega^{\prime})$ in Eqs. (47) and (48) in the hydrodynamic regime and at zero temperature for $y=0.5$ and $k^{\prime}=1.25$ shows different distinct peaks from $z=0-1$ corresponding to the density and the spin waves, centered at $1$ and $\zeta$. Inset: Dependence of $\zeta$ on $z$, illustrating the $p$-wave-induced spin-density separation.
@@
Here, the ground-state compressibility $\chi_{\rho(\sigma)}$ is related to the ground-state energy $E$ through the expression $\chi^{-1}=\frac{1}{V}\frac{\partial^{2}E}{\partial n^{2}}$, where $V$ denotes the constant system volume and $n=N/V$ represents the density. The compressibility for
$\chi_{\rho}^{-1}=g_{\rho}+\frac{2g_{\rho}^{5/2}n_{0}^{1/2}m_{\rho}^{3/2}}{(\pi^{2}\hbar^{3})},$ (42)
and the compressibility for the spin-density degree of freedom is given by
$\chi^{-1}_{\sigma}$ $=g_{\sigma}+\frac{2g_{\sigma}^{5/2}m^{3/2}n_{0}^{1/2}z^{2}}{\pi^{2}\hbar^{3}(z+1)^{7/2}}$
$\quad-\frac{4g_{\sigma}^{5/2}m^{3/2}n_{0}^{1/2}z}{\pi^{2}\hbar^{3}(z+1)^{5/2}}+\frac{2g_{\sigma}^{5/2}m^{3/2}n_{0}^{1/2}}{\pi^{2}\hbar^{3}(z+1)^{3/2}}.$ (43)
The sound velocity is similarly determined as $v_{\rho\left(\sigma\right)}=\left(\frac{V}{mn}\frac{\partial^{2}E}{\partial V^{2}}\right)^{1/2}$ with a constant particle number $N$, expressed for the density degree of freedom as
$v_{\rho}=\left(\frac{g_{\rho}n_{0}}{m_{\rho}}+\frac{2g_{\rho}^{5/2}m_{\rho}^{1/2}n_{0}^{3/2}}{\pi^{2}\hbar^{3}}\right)^{1/2},$ (44)
and for the spin-density degree of freedom as
$v_{\sigma}$ $=\left(\frac{2g_{\sigma}^{5/2}m^{1/2}n_{0}^{3/2}z^{2}}{\pi^{2}\hbar^{3}(z+1)^{5/2}}+\frac{2g_{\sigma}^{5/2}m^{1/2}n_{0}^{3/2}}{\pi^{2}\hbar^{3}(z+1)^{1/2}}\right.$
$\quad\left.-\frac{4g_{\sigma}^{5/2}m^{1/2}n_{0}^{3/2}z}{\pi^{2}\hbar^{3}(z+1)^{3/2}}+\frac{g_{\sigma}n_{0}(1+z)}{m}\right)^{1/2}.$ (45)
Additionally, in calculating the damping rates $\Gamma_{\rho,\sigma}$, we follow Ref. [32] by introducing the Beliaev damping formalism developed for a single-component $s$-wave Bose gas, which determines the width of the DSF peaks. We assume that the $p$-wave interaction effects are primarily captured by modifying the spin-density mode’s effective mass $\left(m_{\sigma}=m/\left(1+z\right)\right)$, while the density mode retains the bare mass $m$, as derived in Eq. (30). Based on this assumption, we approximate the damping rates as $\Gamma_{\rho\left(\sigma\right)}=3\hbar k^{5}/\left(640\pi m_{\rho\left(\sigma\right)}n_{0}\right)$, incorporating the respective effective masses for the density and spin-density modes. When the $p$-wave interaction vanishes $\left(z=0\right)$, the damping rate reduces to the $s$-wave form, consistent with Refs. [32, 62]. This $k^{5}$ dependent form arises from low-energy quasiparticle scattering, with $p$-wave influences indirectly reflected through the mass terms. Importantly, our study focuses on the spin-density separation, driven by the energy splitting between the spin and density modes as given by the excitation energies $\varepsilon_{\rho}$ and $\varepsilon_{\sigma}$ (Eq. (39)), while $\Gamma_{\rho,\sigma}$ affects only the broadening of the DSF peaks without significantly altering this separation. Using these damping rates, we reexpress the DSFs as follows:
$S_{\rho\left(\sigma\right)}\left(\omega\right)$ $=$ $\chi_{\rho}J_{\rho\left(\sigma\right)}\left(\omega^{\prime}\right),$ (46)
@@
$J_{\rho}\left(\omega^{\prime}\right)$ $=$ $\frac{\Gamma_{\rho}/v_{\rho}k^{\prime}}{\left(\omega^{\prime}-1\right)^{2}+\left(\Gamma_{\rho}/v_{\rho}k^{\prime}\right)^{2}},$ (47)
@@
$J_{\sigma}\left(\omega^{\prime}\right)$ $=$ $\frac{\alpha\Gamma_{\sigma}/v_{\rho}k^{\prime}}{\left(\omega^{\prime}-\zeta\right)^{2}+\left(\Gamma_{\sigma}/v_{\rho}k\right)^{2}}.$ (48)
[…]
@@ V CONCLUSION
The primary contribution of this article lies in deriving the EOS under $p$-wave interactions and elucidating their influence on a two-component Bose-Bose mixture through spin-density separation. Beyond providing explicit expressions for the LHY-order corrections to the ground-state energy and quantum depletion, as detailed in Sec. III, this work offers a visualization scheme via the DSFs and a comprehensive analysis of the theoretical results, as presented in Sec. IV. A key finding is that $p$-wave interactions, parameterized by $z$, induce a tunable splitting in the DSFs, reflecting affected particle-number fluctuations in the spin degrees of freedom, which distinguishes this system from purely $s$-wave-dominated mixtures. Furthermore, the validity of the Bogoliubov approximation employed in our calculations is retrospectively confirmed by estimating the quantum depletion $\left(N-N_{0}\right)/N$, utilizing Fig. 2 (b). In typical Bose-Bose mixture experiments, with parameters $n\approx 3\times 10^{14}\leavevmode\nobreak\ \text{cm}^{-3}$, $a\approx 2.7517\leavevmode\nobreak\ \text{nm}$ [63], the quantum depletion in Eq. (26) is evaluated as $\left(N-N_{0}\right)/N\approx 0.0028\times I\left(z\right)$ shown in Fig. 2 (b), affirming the applicability of the Bogoliubov approximation [64].In addition, we briefly address the applicability of our single-channel model in the context of $p$-wave interactions, particularly near Feshbach resonances where a two-channel model might be more appropriate due to the coupling between open and closed channels. Our analysis employs a single-channel framework with the $p$-wave coupling $g_{p}=2\pi\hbar^{2}\nu_{p}/m$, tuned via the scattering volume $\nu_{p}$, which is sufficient to capture the weak-interaction regime ($10<y<$ 1 , $z$ from 00 to $2$) explored in this work. This regime, far from the strong-coupling limit of $p$-wave Feshbach resonances, aligns with experimental conditions where $\nu_{p}$ remains moderate and single-channel descriptions hold, as valid ated by the tunability demonstrated in Ref. [15]. While a two-channel model could provide a more comprehensive description near resonance, enhancing predictions of resonance-enhanced effects, our focus on the Gaussian-level LHY corrections and spin-density separation justifies the simplicity and efficacy of the single-channel approach within the specified parameter space.
The experimental realization of our scenario relies on precise control of three key parameters: the $s$-wave inter- and intra-species scattering lengths ($a_{12}$ and $a$) and the $p$-wave scattering volume ($v_{p}$), all of which are highly tunable via Feshbach resonance [59, 15]. We anticipate that these predicted features, particularly the $p$-wave-induced spin-density separation observable through Bragg spectroscopy, will be validated in future experiments. However, the treatment of Beliaev damping under $p$-wave interactions remains an open question deserving further study. While we have approximated the damping rates using an $s$-wave formalism with effective mass adjustments, a detailed analysis incorporating $p$-wave-specific scattering processes could refine our understanding of the DSF peak broadening and its experimental implications.
@@
In summary, this paper examines the effects of $p$-wave interactions on the EOS and spin-density separation within a Bose-Bose mixture. At the Gaussian order, we demonstrate that $p$-wave interactions fully decouple the spin-density degree of freedom, resulting in a compression effect confined to its effective mass, as evidenced by the renormalization captured in Sec. IV. This decoupling, driven by the momentum-dependent nature of $p$-wave interactions, not only manifests as distinct DSF peaks but also enriches the system’s phase behavior beyond traditional $s$-wave frameworks. However, the treatment of Beliaev damping under $p$-wave interactions remains an open question deserving further study. While we have approximated the damping rates using an $s$-wave formalism with effective mass adjustments, a detailed analysis incorporating $p$-wave-specific scattering processes could refine our understanding of the DSF peak broadening and its experimental implications. By bridging theoretical predictions with experimental feasibility, this study enhances our understanding of interaction-driven phenomena in multi-component Bose systems and provides robust theoretical guidance for future explorations in higher-dimensional quantum gases, potentially advancing the study of exotic quantum states influenced by $p$-wave interactions.
VI ACKNOWLEDGMENTS